OpenStudy (thecalchater):

Given h(x)=2cos x-sin 2x, according to the First Derivative Test for what values of x does h have a relative minimum on the interval (-pi,pi)? x=-5pi/6 x=-5pi/6,-pi/6 x=-5pi/6,-pi/6,pi/2 x=pi/2 What are the points of inflection of k(x)=sin -1/4 sin 2x? x=-pi/3 x=-pi/6 x=-pi/3,pi/x x=-pi/3,0,pi/3

OpenStudy (thecalchater):

For the second one I know it is either C or D.

jigglypuff314 (jigglypuff314):

@Loser66 @welshfella

OpenStudy (welshfella):

First find the derivative of cos x - sin 2x can you do that?

OpenStudy (welshfella):

sorry i have to go right now but I see Loser66 is here I'm sure he'll be able to help

OpenStudy (loser66):

The first derivative of h(x) is ?? your answer above is not correct

OpenStudy (thecalchater):

he gave me a different function of h(x) he dropped the 2 from the front f the cos

OpenStudy (loser66):

type it here, please

OpenStudy (thecalchater):

the derivative of h(x) is -2cos 2x - 2 sin x

OpenStudy (loser66):

I meant the original h(x) =?

OpenStudy (thecalchater):

2 cos x-sin 2x

OpenStudy (thecalchater):

Am I correct for the second question with c or d?

OpenStudy (loser66):

solve one by one, please, the first one, h(x) = 2 cos x - sin(2x), sure?

OpenStudy (thecalchater):

yeah

OpenStudy (loser66):

h'(x) = (2cos(x) )' - (sin(2x))' = -2sin(x) - 2cos (2x) Now, set it =0 and solve for x. What do you get?

OpenStudy (thecalchater):

I dont understand?

OpenStudy (thecalchater):

When I plug it into a calculator it says ERROR

OpenStudy (loser66):

no way to solve by calculator, you must solve it by hand

OpenStudy (thecalchater):

can you walk me through the steps?

OpenStudy (thecalchater):

I'm sorry I never had pre calc

OpenStudy (loser66):

What?? no precal? how to solve cal problem?

OpenStudy (thecalchater):

They made me skip it guidance counselor told me a could handle it and i am barely passing

OpenStudy (loser66):

ok

OpenStudy (thecalchater):

ok so can you walk me through it ?

OpenStudy (loser66):

I don't, but wolfram can help. read it carefully. http://www.wolframalpha.com/input/?i=-2sinx+-+2cos%282x%29+%3D0+

OpenStudy (thecalchater):

So which is the answer? I am leaning towards C

OpenStudy (loser66):

last one, x =pi/2

OpenStudy (thecalchater):

ok thank you now to the second one. I think it is c or d I cant figure out whether 0 is included or not

OpenStudy (thecalchater):

@jim_thompson5910 @pooja195 @TheSmartOne

jimthompson5910 (jim_thompson5910):

The second one has \[\Large k(x) = \sin(x) - \frac{1}{4}\sin(2x)\] right? or no?